OpenStudy (anonymous):

52 people were asked about the number of years they had been in college. The responses are in the table below. Find the expected number of years of college attended. Years of college number attended 0 7 1 6 2 7 3 9 4 9 5 14

OpenStudy (kropot72):

To find the expected number of years of college attended The following formula is used: \[E(X)=\sum_{}^{}xp(x)\]  \[E(X)=(0\times \frac{7}{52})+(1\times \frac{6}{52})+(2\times \frac{7}{52})+.......+(5\times \frac{14}{52})\] Can you now plug in the missing terms and find the value of E(X)?

OpenStudy (kropot72):

@randellcak Are you there?

OpenStudy (anonymous):

Thank you for replying. I have been out of town and unable to check back until today.

OpenStudy (kropot72):

Okay :)